OpenStudy (anonymous):

During the production of sulphuric acid by the Contact process, sulphur dioxide is converted to sulphur trioxide according to the equation: SO2(g) + O2(g) -> 2SO3(g) Δ H = -285.8kJ/mol Calculate the heat energy released in the production of one tonne (1.00x106 g) of sulphur trioxide gas. How do I do a question like this?

OpenStudy (anonymous):

Look at the units. It says KJ per mol. You should be thinking " I love mols and I love grams! I also know how to convert between the two! If I have the amount of energy produced for every mole and I can figure out the moles from the grams I have, then this is \[\Huge{Easy \ peasy \color{yellow}{\ lemon} \ squeasy!}\]

OpenStudy (anonymous):

\[n = \frac{mass}{molecular formula}\] \[n = \frac{106g}{32+3(16)g/mol}\] =1.325 moles 285.8 KJ per mole = so multiply

OpenStudy (anonymous):

@thushananth01 I'm pretty sure he means\[1.00\times 10^6g\]

OpenStudy (anonymous):

(1.00x106 g) he has mentioned it like this...anyway if so he would understand how to substitute the values:)

OpenStudy (anonymous):

One tonne means\[1.00 \times 10^6\] silly goose.